The most recent spacecraft telemetry was acquired on Nov.1 from the Deep Space Network tracking complex at Madrid, Spain. The Cassini spacecraft is in an excellent state of health and with the exception of the CAPS instrument being powered off, all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the "Present Position" page at: http://saturn.jpl.nasa.gov/mission/presentposition/.
Wednesday, Oct. 26 (DOY 299)
An image of Saturn's rings with Titan, Dione, Pandora and Pan is today's Astronomy Picture of the Day. It is available at: http://apod.nasa.gov/apod/ap111026.html. The original post is here: http://saturn.jpl.nasa.gov/photos/imagedetails/index.cfm?imageId=4384.
Thursday, Oct. 27 (DOY 300)
Today the Imaging Science Subsystem (ISS) narrow angle camera had an unplanned warm start during an idle period and recovered nominally. No images were lost, however, incident surprise anomaly (ISA) #50468 was issued.
Friday, Oct. 28 (DOY 301)
Today a member of the Cassini Outreach team led a teacher professional development workshop for 27 teachers from New York, New Jersey, New Hampshire, Connecticut, and Massachusetts at the National Science Teachers' Association regional conference in Hartford, CT. The workshop was entitled "Our Solar System Through the Eyes of Scientists", and included hands-on activities as well as videos, web content, and an introduction to a new NASA solar system thematic curriculum for students in grades 1-6. This curriculum is a follow-on to Cassini's "Reading, Writing & Rings" science and language arts curriculum.
Orbit Trim Maneuver (OTM) #297 was performed today. This was the apoapsis maneuver setting up for the Enceladus 16 encounter on Nov. 6. The Reaction Control Subsystem (RCS) burn began at 5:45AM PDT. Telemetry immediately after the maneuver showed a burn duration of 40.25 seconds, giving a delta-V of 0.046 m/s, and making an 80 km correction in the Enceladus 16 B-plane. All subsystems reported nominal performance after the maneuver. The next maneuver, OTM-298, is scheduled to execute on Wednesday, November 2, and is the final targeting maneuver for E-16. Following OTM298, OTM-299, the post E-16 clean-up maneuver, is scheduled to execute on Tuesday, Nov. 8, with a deterministic delta-V of 2.2 m/s.
Sunday, Oct. 30 (DOY 303)
This week while the spacecraft was near apoapsis, science observations were dominated by Titan observations, with a series of three 13.5 hour and one 4.75 hour observations by the Composite Infrared Spectrometer (CIRS) and Imaging Science Subsystem (ISS). The Cosmic Dust Analyzer (CDA) performed two 15 hour and one 13 hour interstellar dust observations.
Monday, Oct. 31 (DOY 304)
S72 Sequence Implementation Process (SIP) port 3 products were due today. The products will be merged and sent out to the flight team for review.
A kickoff meeting was held today for the S73 Sequence Implementation Process. Port 1 for the first set of input files from the teams occurs on Nov. 14.
Tuesday, Nov. 1 (DOY 305)
The Radio Science (RSS) Team completed Solar Corona Experiment #8; the experiment went well overall. The next RSS data collection occurs on DOY 325.
In support of S73 sequence development, the Attitude and Articulation Control Subsystem (AACS) Team completed its analysis of Titan flybys T-83 and T-84, now that these pointing designs are considered "frozen". This analysis was performed to assess whether or not there are any control authority issues for these upcoming relatively low altitude flybys on thrusters.
The Downlink Ground System (DGS) team delivered Acelog V2.0.1 via a Delivery Coordination Meeting (DCM) held Thursday, Oct. 27, and installed this latest patch release into operations today. The DGS team also installed Mission Control, Data Mgmt, & S/C Analysis (MDAS) V5.0.6 today, to provide an Automatic Alarm Notification (AAN) hysteresis fix.